Use the following information to determine the ROA for the firm in Fiscal 2012

Precision Tool
2012 Income Statement


Net sales                                             $36,408
Less: Cost of goods sold 28,225
Less: Depreciation                                  1,760
Earnings before interest and taxes 6,423
Less: Interest paid 510
Taxable Income $5,913
Less: Taxes 2,070
Net income $3,843


                                                            Precision Tool
                                                2011 and 2012 Balance Sheets


2011             2012 2011               2012
Cash $2,060 $1,003             Accounts payable $7,250 $8,384
Accounts rec. 3,411 4,218             Long-term debt 9,800 11,500
Inventory 18,776 21,908             Common stock 15,000 17,500
Total $24,247 $27,129             Retained earning 6,357 3,825
Net fixed assets 14,160 14,080             Total assets $38,407 $41,209

Total liab & equity $38,407 $41,209
Answer #1

Answer:

Return on Assets (ROA) = Net Income / Average assets

Average asset of Precision Tool in Fiscal 2012 = ($38,407 + $41,209) / 2 =$39,808

Given:

Net Income for Fiscal 2012 = $3,843

Hence:

ROA = 3843 / 39808 *100 = 9.65%

ROA for the firm in Fiscal 2012 = 9.65%